OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. Angle P and Angle Q are supplementary angles. If Angle Q is ten less than four times Angle P, what are the measures of the angles?

OpenStudy (anonymous):

Easy, Q = 4P - 10. We know Q + P = 180. So use substitution, 4P - 10 + P = 180. 5P - 10 = 180. 5P = 190. P = 190/5. Which is 38. So P is 38. Q = 180-38 = 142.